The system contains 1025 molecules of air. Estimate the energy required to raise the system temperature by 5 ° C. Take the Boltzmann constant from the corresponding tables. (K = 1.38 × 10-23)


1
Expert's answer
2021-03-29T09:06:36-0400

Find the amount of substance of 1025 molecules of air assuming this is a diatomic gas:


"n=\frac{N}{N_a}."


The energy required to raise the system temperature by 5 °C is


"Q=\frac 52 nk\Delta T=\frac{5N}{2N_A}k\Delta T,\\\space\\
Q=2.87·10^{-21}\text{ J}."